OpenStudy (anonymous):

Write the equation for the horizontal line passing through (4, -6).

OpenStudy (anonymous):

Horizonatal lines go which way? This can show you the way to go with the equation.

OpenStudy (anonymous):

horizontal line is a line that is parallel to the x axis or the line y=0. since parallel lines have same slope, the slope of this line is 0. \[\frac{y-y _{1}}{x-x_{1}} = m \] \[\frac{y-(-6)}{x-4}=0\] \[y+6=0\] \[y=-6\] is the equation.

OpenStudy (anonymous):

OMG thank you so much!

OpenStudy (anonymous):

Yes, since horizontal lines never touch x-axis their equations are always y= . That's one way to remember.